A triangle has angles of (2x + 11) degrees, (3x - 2) degrees and 4x degrees. Find each angle.
Answer by stanbon(57361)

A triangle has angles of (2x + 11) degrees, (3x - 2) degrees and 4x degrees. Find each angle.
---------------------
The sum of the angles is 180.
2x+11 + 3x-2 + 4x = 180
9x + 9 = 180
x + 1 = 20
x = 19
-----------
2x+11 = 2*19+11 = 49 degrees
3x-2 = 3*19-2 = 55 degrees
4x = 4*19 = 76 degrees
===============
Cheers,
Stan H.